OpenStudy (anonymous):

Solve cos^(2)(x) - 3sin(x) = 3 for x on [0, 2pi] So I have simplified the function down to (-sin(x) - 2)(sin(x) + 1) = 0 or (-sin(x) -1)(sin(x) + 2) = 0 I have determined x = 3pi/2 as sin(3pi/2) = 1/2 Is that the only value possible As I don't think there is a value in the domain of this function that I could use to solve (sin(x) + 2) or (-sin(x) - 2) I could imput arcsin(2) but would that be wrong and if so why?

OpenStudy (watchmath):

because the value of sin is at most 1. SO arcsin(2) is undefined. So the solutions only comes from -sin x -1 =0

OpenStudy (anonymous):

sin(3pi/2) = -1

OpenStudy (anonymous):

made a typo. so I'm correct.

OpenStudy (watchmath):

yes \(e\pi/2\) is the only answer

OpenStudy (anonymous):

so I should always adhere to the range of the unit circle when dealing with trig functions?

OpenStudy (watchmath):

I mean \(3\pi/2\)

OpenStudy (anonymous):

yeah I was wondering lol